# Asymptotic and compound e-values: multiple testing and empirical Bayes

### v3

@@ 4.2 Every FDR procedure is an instance of e-BH
4.3 Combination and derandomization
The connections between FDR controlling procedures, e-BH, and compound e-values laid out in Sections 4.1 and 4.2 motivate the following general and practical mechanism for combining discoveries across multiple testing procedures. To be concrete, suppose we run $L$ different multiple testing procedures $\mathcal{D}_{1},\dots,\mathcal{D}_{L}$. The $\ell$-th procedure is applied to a subset of hypotheses $\{\mathcal{P}_{k}:\;k\in\mathcal{\mathcal{missing}}{S}_{\ell}\}$ where $\mathcal{S}_{\ell}\subseteq\mathcal{K}$ and controls the FDR at level $\alpha_{\ell}$. Then we may proceed as follows:
  1. 1.
    For the $\ell$-th multiple testing procedure, form compound e-values $E_{k}^{(\ell)},\,k\in\mathcal{S}_{\ell}$, for $\mathcal{P}_{k},\,k\in\mathcal{S}_{\ell}$ (which always exist by Theorem 4.2). They could be—but need not necessarily be—the implied ones formed in Theorem 4.2. For $k\in\mathcal{K}\setminus\mathcal{S}_{\ell}$, set $E_{k}^{(\ell)}=1$.
  2. 2.
  3. 3.
The above construction is guaranteed to control the FDR at level $\alpha$, as long as all the individual procedures $\mathcal{D}_{\ell}$ form valid compound e-values. Note that the value of $\alpha_{\ell}$ does not matter, as it is only used in the construction of $E_{1}^{(\ell)},\dots,E_{K}^{(\ell)}$ , possibly implicitly (e.g., in Theorem 4.2).
One important application of the above recipe is derandomization. Suppose that $\mathcal{D}_{\ell}$ is a randomized multiple testing procedure, that is, it is a function of both the data $X$ as well as a random variable $U_{\ell}$ generated during the analysis. Such randomness may not be desirable, since different random number generation seeds will lead to different sets of discoveries. In such cases, the above recipe can be used to construct a new derandomized procedure $\mathcal{D}$ that is less sensitive to $U_{1},\dots,U_{L}$ (formally, full derandomization would occur if $U_{\ell}$ are iid and $L\to\infty$). In this way Ren and Barber (2024) were able to derandomize the model-X knockoff filter (Candès et al., 2018), a flexible set of methods for variable selection in regression with finite-sample FDR control that previously relied on additional randomness.
Some further applications include the following: Banerjee et al. (2023) used the above recipe for meta-analysis in which the $\ell$-th study only reports the set of tested hypotheses, the set of discoveries, as well as the targeted FDR level. Li and Zhang (2023, 2025) considered several applications including the development of new methods for multiple testing with side information.
We end this subsection by using our recipe to derandomize the randomized e-BH procedures (including $R_{1}$-e-BH, $R_{2}$-e-BH, $R$-e-BH, and $U$-e-BH) proposed by Xu and Ramdas (2024). For instance, $U$-e-BH draws $U\sim\mathrm{Unif}[0,1]$ and then applies e-BH to $E_{1}/U,\ldots,E_{K}/U$. Denote any of these procedures by $\mathcal{D}$. All of these procedures have in common the following properties: they operate on e-values $E_{1},\ldots,E_{K}$ as well as an external source of randomization $U$, they control the FDR without requiring further additional assumptions over e-BH, and finally they make at least as many discoveries as e-BH (that is, $\mathcal{D}\supseteq\mathcal{D}_{\text{e-BH}}$ almost surely), often making strictly more discoveries. Both Xu and Ramdas (2024) and Lee and Ren (2024) have noted the challenge of derandomizing these procedures; naive attempts lead back to e-BH applied to the original e-values. Now consider applying the derandomization procedure of this section with $\mathcal{D}_{\ell}$ denoting an application of $\mathcal{D}$ with external randomization seed $U_{\ell}$, $S_{\ell}=\mathcal{K}$, and weights $w_{\ell}=1$. Then the resulting procedure always make at least as many discoveries as e-BH and is not identical to e-BH (i.e., it may make more discoveries).
4.4 Compound e-values as weights in p-value based multiple testing
Ignatiadis et al. (2024) explained that e-values can be used as weights in p-value based multiple testing. We explain why compound e-values are the natural notion of weights for multiple testing with the p-BH procedure. We first define the p-BH procedure of Benjamini and Hochberg (1995) for FDR control based on p-values, as well as weighted generalizations thereof.
Definition 4.6 (p-BH, weighted p-BH, ep-BH).
with the convention $\max(\varnothing)=0$. When $W_{k}=1$ for all $k$, then the above procedure is the p-BH procedure of Benjamini and Hochberg (1995). When $W_{k}=w_{k}$ is deterministic and such that $\sum_{k\in\mathcal{K}}w_{k}=K$, then the above procedure is the weighted p-BH procedure of Genovese et al. (2006). When $W_{k}=E_{k}$ and $E_{1},\dots,E_{K}$ are compound e-variables, then the above procedure is the ep-BH (e-weighted p-BH) procedure of Ignatiadis et al. (2024).
The generalization of p-BH to allow for deterministic weights by Genovese et al. (2006) was motivated by differentially prioritizing hypotheses to improve the power of p-BH while maintaining frequentist FDR control. The requirement $\sum_{k\in\mathcal{K}}w_{k}=K$ may be interpreted as a fixed size budget to be split across hypotheses. When the p-values are positive regression dependent on a subset (PRDS, Benjamini and Yekutieli (2001)) under the distributions in $\mathcal{P}$ and the weights are deterministic then existing arguments in the literature (Ramdas et al., 2019) demonstrate that the FDR of the weighted p-BH procedure is controlled at $(\alpha/K)\sum_{k\in\mathcal{N}}w_{k}$ under this assumption, which is bounded by $\alpha$ under the budget constraint $\sum_{k\in\mathcal{K}}w_{k}=K$. If the weights $W_{1},\dots,W_{K}$ are random and independent of $P_{1},\dots,P_{K}$, then by applying the above argument conditional on the weights and by iterated expectation, it follows that the FDR is controlled at $\sup_{\mathbb{P}\in\mathcal{P}}(\alpha/K)\mathbb{\sum}_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[W_{k}]$. The latter is controlled at $\alpha$ precisely when $W_{1},\dots,W_{K}$ are compound e-values. We note that Ignatiadis et al. (2024) defined the ep-BH procedure and proved FDR control (in their Theorem 4) when $E_{1},\dots,E_{K}$ are e-values. However, Ignatiadis et al. (2024, Section 4.5) explained that the guarantees would continue to hold for compound e-values (although their definition is slightly different from ours).
In Appendix A we also show that compound e-variables can be used as weights for merging p-variables.
4.5 Compound p-values in multiple testing
So far in this section we have focused on the role of compound e-values in multiple testing. Here we briefly discuss some known results about compound p-values. Armstrong (2022) provided a counterexample of jointly independent compound p-values $P_{1},\dots,P_{K}$, such that the p-BH procedure (Definition 4.6) does not control the FDR.
On the other hand, we show the following result, slightly generalizing a claim by Armstrong (2022).
Proposition 4.7.
If $P_{1},\dots,P_{K}$ are (asymptotic) compound p-values, then the p-BY (Benjamini and Yekutieli, 2001) procedure (that is, the p-BH procedure applied at level $\alpha^{\prime}=\alpha\sum_{k\in\mathcal{K}}1/k$) provides (asymptotic) control of the FDR at $\alpha$ under arbitrary dependence among $P_{1},\dots,P_{K}$.
@@ 4.5 Compound p-values in multiple testing
The nonasymptotic result was proved by Armstrong (2022), but here is an alternate proof that makes it transparent to extend to the asymptotic case. Wang and Ramdas (2022) proved that the p-BY procedure may be equivalently described by two steps. First, the p-values $P_{k}$ are calibrated to e-values $E_{k}$ via a specific choice of calibrator $f:p\mapsto T(\alpha/(\ell_{K}p))/\alpha$, where
Second, the e-BH procedure is applied to $E_{1},\dots,E_{K}$. With this, the nonasymptotic claim then follows directly from Theorems 2.3 and 4.4. The asymptotic claim follows because calibrating asymptotic p-values yields asymptotic e-values, which when used with the e-BH procedure provides asymptotic FDR control. ∎
Furthermore, Armstrong (2022) proved that the p-BH procedure with compound p-values controls the FDR under weak dependence asymptotics (as in Storey et al. (2004)); such a property of the p-BH procedure (with compound p-values) was also anticipated earlier by Efron (2007).
@@ 5.2 The optimal discovery compound e-values
$\sum_{k:\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}}\mathbb{E}^{\mathbb{P}}[E_{k}]=\sum_{k:\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}}\int\frac{K\overline{q}^{\circ}(x)}{\sum_{j\in\mathcal{K}}p_{j}^{\circ}(x)}p_{k}^{\circ}(x)\nu(\mathrm{d}x)\leq\int K\overline{q}^{\circ}(x)\nu(\mathrm{d}x)=K,$
@@
The interpretation of the result is as follows (Storey, 2007, Section 5.2). Suppose for each $k\in\mathcal{K}$ we draw $X_{k}$ with probability $\pi_{0}\in(0,1)$ from $\mathbb{P}_{k}$ and probability $1-\pi_{0}$ from $\mathbb{Q}_{k}$. Then the procedure that thresholds the optimal discovery statistics $s(X_{k})$ maximizes the expected number of true discoveries among all simple separable multiple testing procedures with the same expected number of false discoveries. Together with Theorem 5.2, this motivates our terminology: we call $s(X_{1}),\ldots,s(X_{k})$ the optimal discovery compound e-values
@@
  • The objective $\mathbb{E}[\log(s)]$ that we use in (9) is a long-standing criterion of optimality; see the classic works of Kelly (1956); Breiman (1961); Bell and Cover (1980), and the more recent works of Shafer (2021); Grünwald et al. (2024); Vovk and Wang (2024); Waudby-Smith and Ramdas (2024). This quantity represents the rate of growth of e-processes in a single-hypothesis setting. Other choices of the objective can be used for (9) and they would lead to different forms of the optimal compound e-values.
  • The objective value of the optimization problem is given by $\operatorname{KL}(\overline{\mathbb{Q}}^{\circ}\,||\,\overline{\mathbb{P}}^{\circ})$, where $\operatorname{KL}(\cdot\,||\,\cdot)$ denotes the Kullback-Leibler divergence. By convexity, $\operatorname{KL}(\overline{\mathbb{Q}}^{\circ}\,||\,\overline{\mathbb{P}}^{\circ})\leq\sum_{k\in\mathcal{K}}\operatorname{KL}(\mathbb{Q}_{k}^{\circ}\,||\,\mathbb{P}_{k}^{\circ})/K$, which means that the optimal discovery compound e-values have a worse objective than using the optimal separable e-value for each individual testing problem of $\mathbb{P}_{k}^{\circ}$ vs $\mathbb{Q}_{k}^{\circ}$ given by $E_{k}^{*}=q_{k}^{\circ}(X_{k})/p_{k}^{\circ}(X_{k})$. In view of the above, why are the optimal discovery compound e-values of interest? The reason is that if $K$ is large (but the individual statistical problems have limited sample sizes), then it becomes possible to mimic the optimal discovery compound e-values without knowledge of the null $\mathbb{P}_{k}^{\circ}$, e.g., via empirical Bayes methods. Such a strategy is fruitful when the hypotheses $\mathcal{P}_{k}$ are composite. We demonstrate such a strategy explicitly in Section 6 on simultaneous t-tests where the nuisance parameters are the variances.
We finally note that any tight simple separable compound e-value for testing point nulls $\mathcal{P}_{k}=\{\mathbb{P}\,:\,\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}\}$ must be of the form (10) for some densities $q_{1},\dots,q_{K}$.
@@
Suppose that $E_{1},\dots,E_{K}$ are tight simple separable compound e-values (with the nulls defined as in the previous parts of this section). Then there exist $\mathrm{d}\nu$-densities $q_{1},\dots,q_{K}$ such that $E_{k}$ may be represented as in (10) on the support of $\sum_{k\in\mathcal{K}}\mathbb{P}_{k}/K$ and it is without loss of generality to take $q_{1}=\dots=q_{K}$.
@@ 5.3 Localized and compound universal inference
We provide an explicit example of such a construction in Section 6.1.1. The second idea is that we can profile over the nuisance parameters appearing in the optimal discovery compound e-values of Theorem 5.2 by lifting to the space of distributions. Compound universal inference (CUI) e-values are defined as follows:666 Nguyen and Gupta (2023) use universal inference for an empirical Bayes application. By contrast, here we use ideas from empirical Bayes and universal inference to provide a new construction of compound e-values.
@@
Suppose that $\widehat{G}(\delta)$ is a $(1-\delta)$-localization for $G(\boldsymbol{\psi})$ for all $\mathbb{P}\in\mathcal{P}$. Then $E_{1}^{\mathrm{CUI}},\ldots,E_{1}^{\mathrm{CUI}}$ defined in (13) are $(0,\delta)$-approximate e-variables for $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$ under $\mathcal{P}$.
@@ 6 The simultaneous t-test sequence model
In this section we provide concrete examples for data-driven implementations of localized compound universal inference, optimal discovery compound e-variables, as well as of asymptotic e-variables and asymptotic compound e-variables under asymptotics in which either the number of hypotheses or the sample size grows (or both). We do so in the stylized but practically relevant setting of multiple testing with simultaneous t-tests. We refer to Smyth (2004); Westfall et al. (2004); Finos and Salmaso (2007); Bourgon et al. (2010); Lu and Stephens (2016); Guo and Romano (2017); Ignatiadis and Huber (2021); Hoff (2022) for further works that have used this setting to demonstrate the fundamental difference between single hypothesis testing and multiple testing.
@@ 6.1 Asymptotic compound e-values under normality
Our setting is encompassed by the composite null setting with nuisance parameters that we described in Section 5.3. The nuisance parameters under the null ($\psi_{k}$ in Section 5.3) are given by the variances $\sigma_{k}^{2}$. Letting $\boldsymbol{\sigma}^{2}=(\sigma_{1}^{2},\ldots,\sigma_{K}^{2})$m=, a central object is the empirical distribution of the $\sigma_{k}^{2}$, that is,
@@ 6.1.1 Localized compound universal inference for the simultaneous t-test
$f_{G}(\hat{\sigma}_{k}^{2})=\int_{0}^{\infty}p(\hat{\sigma}_{k}^{2}\mid\sigma^{2})\mathrm{d}G(\sigma^{2}),\;\;\;\,p(\hat{\sigma}_{k}^{2}\mid\sigma^{2})=\frac{\nu^{\nu/2}}{\left(\sigma^{2}\right)^{\nu/2}2^{\nu/2}\Gamma(\nu/2)}\left(\hat{\sigma}_{k}^{2}\right)^{\nu/2-1}\exp\left(-\frac{\nu\hat{\sigma}_{k}^{2}}{2\sigma^{2}}\right).$
@@
Remark 6.2 (Computation of CUI.).
Computing CUI amounts to solving the optimization problem in (13). To do this in practice, we proceed as follows. First, following Ignatiadis and Wager (2022), we discretize $\mathcal{G}$ by taking a covering of $[\underaccent{\bar}{\sigma}^{2},\bar{\sigma}^{2}]$ with $B$ points $\tilde{\sigma}_{1}^{2},\dots,\tilde{\sigma}_{B}^{2}\in[\underaccent{\bar}{\sigma}^{2},\bar{\sigma}^{2}]$ and considering all distributions of the form $G=\sum_{\ell=1}^{B}\pi_{\ell}\delta_{\tilde{\sigma}_{\ell}^{2}}$, where $(\pi_{1},\dots,\pi_{B})$ lies on the probability simplex. Second, we relax the Kolmogorov-Smirnov constraint in (19) by enforcing it only on a finite subset $\mathcal{T}\subseteq(0,\infty)$; e.g., we can let $\mathcal{T}$ be a subset of the realized sample variances $\hat{\sigma}_{1}^{2},\dotsc,\hat{\sigma}_{K}^{2}$. This second discretization step is always conservative, that is, it makes the localization strictly larger. Our computational approximation to $E_{k}^{\mathrm{CUI}}$ is then given by $q^{\circ}(X_{k})/\mathrm{obj}$, where $\mathrm{obj}$ is the optimal value of the following linear program,

### v4

@@ 4 Compound e-values are central to multiple testing
We first construct compound e-values from $\mathcal{D}$ following Banerjee et al. (2023). Let
$E_{k}=\frac{K}{\alpha}\frac{V_{k}}{R_{\mathcal{D}}\lor 1}.$ (4)
Fix any $\mathbb{P}\in\mathcal{P}$. Then,
where the inequality follows from the definition of FDR guarantee. Thus, $E_{1},\dots,E_{K}$ are indeed compound e-variables for $\mathcal{P}_{1},\dotsc,\mathcal{P}_{K}$ under $\mathcal{P}$,
By Theorem 4.3, e-BH applied to $E_{1},\ldots,E_{K}$ controls the FDR. Moreover, $k_{e}^{*}$ in Definition 4.2 is seen to be equal to $R_{\mathcal{D}}$. Thus, $\mathcal{D}$ and e-BH on $E_{1},\ldots,E_{K}$ make the same discoveries. We postpone the proof of tightness to Appendix G. ∎
Such a result is not known to hold for the BH procedure, or for any other procedure. The e-BH procedure, coupled with the notion of compound e-values, truly stands out in that respect.
Theorem 4.4 captures the fundamental role of compound e-values for multiple testing with FDR control. There are several further practical applications in which compound e-values play a fundamental role in multiple testing. We review four such applications in Appendix C:
  • Appendix C.1 explains how compound e-values enable combination and derandomization of multiple testing procedures, including a new application to derandomization of the randomized e-BH procedures of Xu and Ramdas (2024).
  • Appendix C.2 demonstrates that compound e-values provide a natural notion of weights for the ep-BH procedure of Ignatiadis et al. (2024).
  • Appendix C.3 shows that compound e-values can be used as weights for merging p-values. These results are new.
  • Appendix C.4 briefly describes the role of compound p-values in multiple testing. Compound p-values can be used for multiple testing via calibration to compound e-values and applying the e-BH procedure; this argument generalizes a result of Armstrong (2022) and provides an alternative proof that the Benjamini-Yekutieli procedure applied to compound p-values controls the FDR.
@@ 5.2 The optimal discovery compound e-values
$\sum_{k:\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}}\mathbb{E}^{\mathbb{P}}[s(X_{k})]=\sum_{k:\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}}\int\frac{K\overline{q}^{\circ}(x)}{\sum_{j\in\mathcal{K}}p_{j}^{\circ}(x)}p_{k}^{\circ}(x)\nu(\mathrm{d}x)\leq\int K\overline{q}^{\circ}(x)\nu(\mathrm{d}x)=K,$
@@
The interpretation of the result is as follows (Storey, 2007, Section 5.2). Suppose for each $k\in\mathcal{K}$ we draw $X_{k}$ with probability $\pi_{0}\in(0,1)$ from $\mathbb{P}_{k}$ and probability $1-\pi_{0}$ from $\mathbb{Q}_{k}$. Then the procedure that thresholds the optimal discovery statistics $s(X_{k})$ maximizes the expected number of true discoveries among all simple separable multiple testing procedures with the same expected number of false discoveries. Together with Theorem 5.2, this motivates our terminology: we call $s(X_{1}),\ldots,s(X_{k})$ the optimal discovery compound e-values.
@@
  • The objective $\mathbb{E}[\log(s)]$ that we use in (9) is a long-standing criterion of optimality; see the classic works of Kelly (1956); Breiman (1961); Bell and Cover (1980), and the more recent works of Shafer (2021); Grünwald et al. (2024); Vovk and Wang (2024); Waudby-Smith and Ramdas (2024); Larsson et al. (2025). This quantity represents the rate of growth of e-processes in a single-hypothesis setting. Other utility functions beyond logarithmic utility can be used in (9) and lead to different forms of optimal simple separable compound e-values. Our framework provides a general method for translating single e-value results to the compound setting. In Appendix D, we illustrate this approach using results of Koning (2025) on optimal e-values with more general utility functions.
  • The objective value of the optimization problem is given by $\operatorname{KL}(\overline{\mathbb{Q}}^{\circ}\,||\,\overline{\mathbb{P}}^{\circ})$, where $\operatorname{KL}(\cdot\,||\,\cdot)$ is the Kullback-Leibler divergence. By convexity, $\operatorname{KL}(\overline{\mathbb{Q}}^{\circ}\,||\,\overline{\mathbb{P}}^{\circ})\leq\sum_{k\in\mathcal{K}}\operatorname{KL}(\mathbb{Q}_{k}^{\circ}\,||\,\mathbb{P}_{k}^{\circ})/K$, which means that the optimal discovery compound e-values have a worse objective than using the optimal separable e-value for each individual testing problem of $\mathbb{P}_{k}^{\circ}$ vs $\mathbb{Q}_{k}^{\circ}$ given by $E_{k}^{*}=q_{k}^{\circ}(X_{k})/p_{k}^{\circ}(X_{k})$. In view of the above, why are the optimal discovery compound e-values of interest? The reason is that if $K$ is large (but the individual statistical problems have limited sample sizes), then it becomes possible to mimic the optimal discovery compound e-values without exact knowledge of the null $\mathbb{P}_{k}^{\circ}$, e.g., via empirical Bayes methods. Such a strategy is fruitful when the hypotheses $\mathcal{P}_{k}$ are in fact composite. We demonstrate such a strategy explicitly in Section 6 on simultaneous t-tests where the nuisance parameters are the variances.
We finally note that any tight simple separable compound e-value for testing point nulls $\mathcal{P}_{k}=\{\mathbb{P}\,:\,\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}\}$ must be of the form (10) for some densities $q_{1}^{\circ},\dots,q_{K}^{\circ}$.
@@
Suppose that $E_{1},\dots,E_{K}$ are tight simple separable compound e-values (with the nulls defined as in the previous parts of this section). Then there exist $\mathrm{d}\nu$-densities $q_{1}^{\circ},\dots,q_{K}^{\circ}$ such that $E_{k}$ may be represented as in (10) on the support of $\sum_{k\in\mathcal{K}}\mathbb{P}_{k}^{\circ}/K$ and it is without loss of generality to take $q_{1}^{\circ}=\dots=q_{K}^{\circ}$.
@@ 5.3 Localized and compound universal inference
We provide an explicit example of such a construction in Section 6.1.1. The second idea is that we can profile over the nuisance parameters appearing in the optimal discovery compound e-values of Theorem 5.2 by lifting to the space of distributions. Compound universal inference (CUI) e-values are defined as follows:888 Nguyen and Gupta (2023) use universal inference for an empirical Bayes application. By contrast, here we use ideas from empirical Bayes and universal inference to provide a new construction of compound e-values.
@@
Suppose that $\widehat{G}(\delta)$ is a $(1-\delta)$-localization for $G(\boldsymbol{\psi})$ for all $\mathbb{P}\in\mathcal{P}$. Then $E_{1}^{\mathrm{CUI}},\ldots,E_{1}^{\mathrm{CUI}}$ defined in (13) are $(0,\delta)$-approximate compound e-variables for $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$ under $\mathcal{P}$.
@@ 6 The simultaneous t-test sequence model
In this section we provide concrete examples for data-driven implementations of localized compound universal inference, optimal discovery compound e-variables, as well as of asymptotic e-variables and asymptotic compound e-variables under asymptotics in which either the number of hypotheses or the sample size grows (or both). We do so in the stylized but practically relevant setting of multiple testing with simultaneous t-tests. We refer to Smyth (2004); Westfall et al. (2004); Finos and Salmaso (2007); Bourgon et al. (2010); Lu and Stephens (2016); Guo and Romano (2017); Ignatiadis and Huber (2021); Hoff (2022); Ignatiadis and Sen (2025) for further works that have used this setting to demonstrate the fundamental difference between single hypothesis testing and multiple testing.
@@ 6.1 Asymptotic compound e-values under normality
Our setting is encompassed by the composite null setting with nuisance parameters that we described in Section 5.3. The nuisance parameters under the null ($\psi_{k}$ in Section 5.3) are given by the variances $\sigma_{k}^{2}$. Letting $\boldsymbol{\sigma}^{2}=(\sigma_{1}^{2},\ldots,\sigma_{K}^{2})$, a central object is the empirical distribution of the $\sigma_{k}^{2}$, that is,
@@ 6.1.1 Localized compound universal inference for the simultaneous t-test
$f_{G}(\hat{\sigma}_{k}^{2})=\int_{0}^{\infty}p(\hat{\sigma}_{k}^{2}\mid\sigma^{2})\mathrm{d}G(\sigma^{2}),\;\;\;\,p(\hat{\sigma}_{k}^{2}\mid\sigma^{2})=\frac{\nu^{\nu/2}\left(\hat{\sigma}_{k}^{2}\right)^{\nu/2-1}}{\left(\sigma^{2}\right)^{\nu/2}2^{\nu/2}\Gamma(\nu/2)}\exp\left(-\frac{\nu\hat{\sigma}_{k}^{2}}{2\sigma^{2}}\right),$ (19)
@@
Remark 6.2 (Computation of CUI).
Computing CUI amounts to solving the optimization problem in (13). To do this in practice, we proceed as follows. First, following Ignatiadis and Wager (2022), we discretize $\mathcal{G}$ by taking a covering of $[\underaccent{\bar}{\sigma}^{2},\bar{\sigma}^{2}]$ with $B$ points $\tilde{\sigma}_{1}^{2},\dots,\tilde{\sigma}_{B}^{2}\in[\underaccent{\bar}{\sigma}^{2},\bar{\sigma}^{2}]$ and considering all distributions of the form $G=\sum_{\ell=1}^{B}\pi_{\ell}\delta_{\tilde{\sigma}_{\ell}^{2}}$, where $(\pi_{1},\dots,\pi_{B})$ lies on the probability simplex. Second, we relax the Kolmogorov-Smirnov constraint in (20) by enforcing it only on a finite subset $\mathcal{T}\subseteq(0,\infty)$; e.g., we can let $\mathcal{T}$ be a subset of the realized sample variances $\hat{\sigma}_{1}^{2},\dotsc,\hat{\sigma}_{K}^{2}$. This second discretization step is always conservative, that is, it makes the localization strictly larger. Our computational approximation to $E_{k}^{\mathrm{CUI}}$ is then given by $q^{\circ}(X_{k})/\mathrm{obj}$, where $\mathrm{obj}$ is the optimal value of the following linear program: